Journalist Report – December 24th

Journalist Report

Ben Durkee, Crew 218 Journalist

Sol 02

The meteorological crystal ball predicted some nasty weather for today, so we were unauthorized to have any EVA’s to avoid being caught in it. On the bright side, this meant we were able to sleep in a bit today.

We metamorphosed from our sleeping bag cocoons toasty and well-rested this morning. Since we were to be cooped up in the Hab all day, we began our morning with some yoga and breathing exercises to stay in shape. After following our rigid Excel spreadsheet of lethargic calisthenics, we tackled our day in a meditative state.

This zen didn’t last long, though, as the heavens promptly opened and proved the forecast right. Apparently, when it rains on Mars, it pours (in a solid-state). Albeit destructive to our scientific progress, we couldn’t help but marvel at the beauty of a Martian snowfall.

Since we had already warmed up our bodies, it was time to sharpen our minds. We whipped out a puzzle to pass the time and may have bitten off more than we could chew. This Mars-themed jigsaw has proven very difficult, what with it being 60% non-descript black space. After a while we all started getting lost in the void of ebony puzzle pieces.

We only have two personal research projects that can be performed without EVA, so LuzMa and Shefali peeled off to work on those as we decided we needed a break from the puzzle. I’m genuinely unsure and impressed at how they had the mental fortitude to be productive after all that – at that point the puzzle had sapped me of all my (admittedly scarce) brain power. Unfortunately Shefali’s research met an untimely roadblock due to computer issues, but I gather that LuzMa’s work is going well!

As I write this, we are finishing dinner and preparing to dive back into the mind-bending cardboard abyss. The snow is slowing, but it has accumulated more than enough to keep us in here for another Sol at a minimum. Though we are confined to the Hab tomorrow, we can only embrace the variables out of our control and enjoy a lovely White Christmars on the red planet.

I’m sure the audience on Earth are waiting with bated breath to hear how the puzzle saga continues, so I will be sure to keep you posted. Consider it a Christmars gift from me to you.

Happy holidays!

Commander Report – December 24th

Happy holidays to all people on old good Earth!

This year I am having a brand new experience here at MDRS. We had heavy snow precipitation today. Snow is not a first for me here on the polar cap of Mars, but the amount is something that I have never seen, and that caused us to be stuck in the habitat. This is a great new experience, getting to see the crew at a more relaxed pace after waking up a little later, to the notes of the song “Vienna”, picked by Ben. Following a Gemini, Apollo, and Space Shuttle tradition, still going on onboard the ISS, I use the crew’s favorite songs as a wake-up call!

After beginning the day with yoga and a pancake-based breakfast (yes, dehydrated eggs and milk worked perfectly), I introduced the crew to the geology of MDRS, and then we had a relaxed time building a puzzle and beginning a research project.

Meanwhile, the snow became to fall on our campus, and the hearts are warmed by a white Christmas which we will have here in isolation, far from our friends and families, but proud to be part of the endeavor to support the exploration of Mars.

A nice day awaits us tomorrow. Ad Astra.

Cesare Guariniello, Commander

Greenhab Report – December 24th

Crew 218 GreenHab Report 24-DEC-19

Crew GreenHab Officer: Dr. Jonathan R. Buzan

Environmental control: Heating.

Shade cloth (40% and 30%) on.

Average temperature: 21.4°C; 13.5% (because of the minimum humidity, I believe this calculation is an error).

10:00AM

Floor Unit: 16°C
Electronic: 23.5°C
humidity 19%

1:55PM

Floor Unit: 15°C
Electronic: 20.4°C
humidity 20%

5:40PM

Floor Unit: 16°C
Electronic: 21.1°C
humidity 20%

Max: 24.9°C; 25%
Min: 17.9°C; 2% (I believe this humidity reading is in error)

Hours of supplemental light: Light system 7:20PM-12:20AM (timer was modified by staff. Last night’s measurement was an estimated start time).

Daily water usage for crops: 5.4Gal.
Daily water usage for research and/or other purposes: N/A

Water in Blue Tank – ~123 Gal.

Time(s) of watering for crops: 5:40 PM

SEASONAL PLANTING HISTORY:

Change to crops: None.

OTHER NOTES:

1st sprouts: Zucchini, Cherry tomato, Buttercrunch.

—No Martian rodents in the trap.

—ALARM: At 2:20 PM an alarm was heard while Shefali and I were in GreenHab. We radioed Habitat to notify of alarm and investigated. Astronomy Observatory’s alarm was set off. No fire detected.

—Used MiracleGrow™ while watering the plants.

Harvest:

—Cilantro plants had coriander seeds. Harvested 0.22g (took seeds to Science Lab to use a triple beam scale).

—Because of limited comms, we warmly ask CAPCOM to search for a dessert (not desert!) recipe that is coriander forward. 😀

Support/supplies requested: N/A

Operations Report – December 24th

Crew 218 Operations Report 24-Dec-19
SOL: 02

Name of person filing report: Luz Maria Agudelo Urrego
Non-nominal systems: NA

Generator: run
Hours run: 14
From what time last night: 1800
To what time this morning: 0800
List any additional daytime hours when the generator was run: N/A

Solar— SOC % (Before the generator is run at night: 79 %

Diesel Reading – 86%
Station Propane Reading – 80%
Ethanol Free Gasoline: N/A

Water (loft tank) (gal): 28
Water Meter (units): 0146980.3
Water (static tank) (gal): 480
Static to Loft Pump used – Yes
Water in Green Hab (gal): 124
Water in Science Dome (gal): 0

Toilet tank emptied: No

Deimos rover used: Still in the workshop
Hours:
Beginning charge:
Ending charge:
Currently charging:

Sojourner rover used: Assigned to director
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: Not used
Hours: 119.2
Beginning charge: (Before EVA):
Ending charge: (On return from EVA, before recharging):
Currently charging: Yes

Opportunity rover used: Still in the workshop
Hours:
Beginning charge:
Ending charge:
Currently charging:

Curiosity rover used: Not used
Hours: 124.4
Beginning charge:
Ending charge: (On return from EVA, before charging):
Currently charging: Yes

Notes on rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No
Reason for use: N/A
Oil Added? No
# Hours the ATVs were used today: 0
Notes on ATVs: N/A

HabCar used and why, where? No
CrewCar used and why, where? No

General notes and comments: N/A

Summary of internet: Covered in snow, Outpost crew cleaned it. Now it is working as usual.
Summary of suits and radios: Nominal
Summary of Hab operations: Nominal
Summary of GreenHab operations: routine watering
Summary of Science Dome operations: Nominal
Summary of RAM operations: Nominal
Summary of any observatory issues: Nominal
Summary of health and safety issues: Nominal

Questions, concerns, and requests to Mission Support: NA

Luz Ma
ENG
Crew 218

Sol Summary – December 24th

Sol: 02

Summary Title: The Weather on Mars is Frightful, but the Hab is so Delightful….

Author’s name: Pat Pesa

Mission Status: Crew is disappointed in the canceling of EVAs, but enjoying the beauty of the snow.

Sol Activity Summary: Crew started the day with some yoga, started some measurements for indoor heat stress analysis, and made significant progress towards a Martian Puzzle.

Look ahead plan: Tomorrow we will stay inside the Hab once more and have some Christmas festivities.

Anomalies in work: None

Weather: Cloudy and Snowing, around 34 degrees

Crew physical status: healthy

EVA: none

Reports to be filed: sol summary, commander report, operations report, greenhab report, journalist report, EVA report

Support Requested: none

Pat
Geologist, MDRS 218